# A Neurosymbolic Framework for Interpretable Cognitive Attack Detection in Augmented Reality

### v1

@@ Appendix A Extended AR-VIM Dataset Samples
Representative sample videos for each attack class (excluding the No Attack category) is shown below. The full dataset consists of 160 videos across diverse scenes.
@@
Figure 6: Sample dataset content for each AR Attack Class.
Appendix B Dataset Construction
For the real-world scenes in the Extended AR-VIM dataset, we collected additional samples using a Meta Quest 3 headset. The data collection setup is shown below
@@ Appendix B Dataset Construction
Appendix C Model Architectures and Training
For our supervised-learning baselines—ResNet and ViViT. Training hyper-parameters are shown below. We conducted all training and evaluation using Google Colab with NVIDIA A100 GPUs (40GB memory). The experiments were run in a Linux environment via Colab’s remote infrastructure, conducted using PyTorch 2.6.0 with Python 3.11.
@@ Appendix C Model Architectures and Training
We use ViViT-B/16x2 (pretrained on Kinetics-400) as the backbone for our ViViT model, and a 3D ResNet-18 for the ResNet baseline. Both models are trained on the full training set using inverse-frequency sampling to balance class distributions. The classification task involves five classes: the four attack types, along with an additional No Attack category. All input videos are resized to 400×400 pixels and normalized using a mean of [0.45, 0.45, 0.45] and a standard deviation of [0.225, 0.225, 0.225]. Each input consists of eight uniformly sampled frames from a 13-second clip. We use Focal Loss with a gamma of 2.0 as the training objective. Mixed precision (FP16) training is enabled via HuggingFace Accelerate for both models. ViViT is trained with a cosine learning rate scheduler, while the 3D ResNet uses a StepLR scheduler (step size = 10, gamma = 0.1). Final hyperparameter configurations for both models were selected based on the settings that yielded the highest validation accuracy.
Appendix D Benchmark Prompt Templates (GPT-o3 and Gemini 2.5)
For our baseline VLM models, we use the following prompt for both GPT-o3 and Google Gemini 2.5. Related work include(Zhang et al. 2025; Zou et al. 2024; Qiao et al. 2024; Zhang et al. 2024b; Jing et al. 2024; Gao et al. 2024; Ravari et al. 2024; Zhang, Aggarwal, and Lan 2025; Yu et al. ; Zhang, Imani, and Lan 2024)
@@
You are an AR video attack analyst. Your job is to identify the presence of 5 different kinds of AR attacks in the video. The definitions for each attack are as seen below: A text modification attack manipulates textual information embedded in the environment, thereby changing the semantic message conveyed by the scene. For example, replacing a NO PARKING FIRE LANE sign with one reading FREE PARKING FIRE LANE, for instance, directly reverses the intended directive and can create significant safety hazards. Such attacks are especially consequential in contexts where text offers primary contextual or navigational cues.
@@
A visual modification attack alters an object’s appearance—its color, shape, or iconography—or repositions it in an implausible manner. Appearance changes, such as transforming a pedestrian-crossing sign into a bicycle-crossing sign or switching a green traffic signal to red, degrade recognition accuracy and mislead both human users and perception systems. Spatial changes include shifting an object away from its logical location or removing it entirely; for example, relocating a stop sign to the centre of an intersection. These manipulations can be quantified by analysing pixel-level discrepancies and deviations from expected spatial distributions.
@@
A removal attack hides or deletes critical objects, thereby eliminating essential information and distorting scene semantics. In a fire evacuation scenario, a blank patch that obscures an emergency-exit sign on a door. By erasing this safety-critical cue, the attack deprives occupants of clear evacuation guidance, potentially endangering lives during an emergency. By analyzing nodes from the scene graph, removal attacks disrupt relational structure and semantic integrity.
@@
An addition attack introduces new, contextually inconsistent, or misleading objects into the scene, thereby distorting its intended semantics or altering user perception. For example, a fake road sign indicating a detour is inserted into a clear street, potentially redirecting traffic and causing confusion. These attacks may include adding fictitious warning signs, virtual obstructions, or irrelevant icons that do not exist in the physical environment. Such manipulations exploit the overlay nature of AR systems and are particularly dangerous when added elements resemble legitimate cues, misleading both human users and perception models. Detection relies on identifying unexpected objects that conflict with the scene’s established structure and context.
@@
The following input will contain a video. Please output the attack kind. Note that there may be more than 1 kind of attack present in the video, or a video could have no attack present at all. Along with the attack, mention the object the attack was performed on or describe the attack briefly in the context of the video.

### v3

@@ Appendix B CADAR-50K Dataset Samples
Representative sample videos for each class (excluding the No Attack category) is shown below. The full dataset consists of 375 videos across diverse scenes.
@@
Figure 8: Sample dataset content for each AR Attack Class. For video examples, please refer to CADAR_demo.mp4 in the supplementary materials.
B.1 Dataset Construction
For the real-world scenes in the CADAR-50K dataset, we collected additional samples using a Meta Quest 3 headset and a cellphone camera. The video dataset contains both malicious cognitive attacks and benign augmented reality content. All data was recorded using the built-in pass-through camera of a Meta Quest headset and a Samsung Galaxy S24. We employed two distinct data generation pipelines to ensure diversity in environmental contexts and lighting conditions.
B.2 In-Situ Augmented Reality Overlay
The first method involved augmenting real-world environments with manipulated overlays to simulate attacks in situ. We first identified target objects within a physical scene (e.g., a store sign displaying a discount). A high-resolution photograph of the target was captured and subjected to digital manipulation. For instance, altering text from ”50% off” to ”150% off” to simulate a text manipulation attack.
The manipulated asset was imported into the Unity engine and mapped onto a planar mesh (Quad). We utilized Unity’s scaling and transformation tools to align the virtual overlay with the physical object’s dimensions and perspective in the real world. To introduce benign confounders, we analyze the scene context and suggest appropriate non-malicious objects. Based on these suggestions, synthetic 3D or 2D assets were generated and anchored within the scene. During the recording phase, a controller input triggered the visibility of the overlay, allowing us to capture the transition between the benign state and the attack state dynamically.
B.3 Synthetic Scene Projection
To expand the dataset beyond accessible physical locations, we utilized a fully synthetic generation pipeline. In the Unity environment, we instantiated a large-scale Quad (approximately $5m\times 5m$) to serve as a dynamic backdrop.
We utilized a generative image model to synthesize realistic background scenes. The generation of scenes containing specific target objects is required for occlusion or manipulation attacks. Once the generated scene was applied to the backdrop, attack artifacts were positioned in the foreground. Although the background was two-dimensional, we achieved a pseudo-three-dimensional effect by strategically angling the foreground objects and the camera perspective. Similar to the in-situ method, the attack artifacts were toggled via controller input during recording to capture paired data samples.
@@ B.3 Synthetic Scene Projection
Appendix C Video Baseline Implementation Details
We provide the full implementation details of the supervised video baselines used in our experiments: the 3D–ResNet model [hara2017learningspatiotemporalfeatures3d] and the transformer-based ViViT model [arnab2021vivitvideovisiontransformer]. Both baselines follow standard practice in video classification and are trained under consistent data splits, preprocessing steps, and evaluation metrics. The descriptions below summarize the implementation in our released training code.
C.1 Dataset Splits and Sampling
All video models use the same dataset partitions:
  • Train / Validation / Test: 60% / 20% / 20%, stratified by attack type and scene category.
  • One clip is extracted per video for all splits, matching the single-view evaluation protocol.
To mitigate class imbalance during training, we adopt a class-balanced sampling strategy. Let $c_{i}$ be the number of training videos for class $i$; class weights are:
$w_{i}=\frac{1}{c_{i}+10^{-6}},\qquad p_{i}=\frac{w_{i}}{\sum_{j}w_{j}},$ (26)
and training videos are sampled according to $p_{i}$. This sampling procedure matches the implementation in our 3D–ResNet code.
C.2 Clip Construction and Preprocessing
For both ViViT and 3D–ResNet, AR videos are processed into short clips using a uniform temporal sampling pipeline:
  • 20-frame clips:
    obtained usingUniformTemporalSubsample(20).
  • Spatial resizing: frames are resized to 400 $\times$ 400.
  • Pixel normalization: mean [0.45,0.45, 0.45], std [0.225, 0.225, 0.225].
Frames are arranged as tensors of shape $(T,C,H,W)$ with $T=20$. No additional augmentations (cropping, flipping, jittering) are applied, avoiding synthetic distortions that could alter AR semantics.
C.3 3D–ResNet Baseline
Our 3D CNN baseline is implemented using the r3d_18 architecture from torchvision (18-layer 3D–ResNet).We detail the full setup below.
Architecture.
  • Model: R3D-18 (3D ResNet).
  • Pretraining: Kinetics-400 weights from Torchvision.
  • Final classifier layer replaced with a 5-way linear head.
Loss Function.
We employ a Focal Loss with $\gamma=2.0$:
$\mathcal{L}_{\mathrm{focal}}=(1-\exp(-\mathrm{CE}(\mathbf{x},y)))^{\gamma}\,\mathrm{CE}(\mathbf{x},y),$ (27)
which improves stability under severe class imbalance.
Optimization.
  • Optimizer: AdamW
[…]
  • Scheduler: StepLR (step=10, $\gamma$=0.1)
  • Batch size: 20
  • Epochs: 20
Precision Training.
Mixed-precision (FP16) training is performed using HuggingFace Accelerate with autocast() enabled, matching our code.All layers are trainable; no freezing is applied.
C.4 ViViT Baseline
We adopt the ViViT-B/16x2 (Factorised Encoder) model as our transformer baseline:
  • Patch size: $\times16$
  • Hidden dimension: 768
  • Transformer layers: 12
  • Heads: 12
  • Pretraining: Kinetics-400
The model receives the same 20-frame clips and performs temporal factorization before spatial tokenization as in the original ViViT design.
Training.
We train ViViT for 20 epochs using:
  • Batch size: 16
  • AdamW optimizer (1e-4 learning rate, 5e-3 weight decay)
  • Cosine learning-rate decay
  • Warmup steps: 10% total steps
  • Mixed precision (FP16) using Accelerate
C.5 Evaluation Metrics
To ensure consistency with the metrics defined in the main paper, we evaluate both video models using the following measures:
Class-Specific Accuracy.
For each class $k$:
$\text{Acc}_{k}=\frac{\mathrm{TP}_{k}}{\mathrm{TP}_{k}+\mathrm{FN}_{k}+\mathrm{FP}_{k}}.$
Precision, Recall, F1.
Computed using scikit-learn with zero_division=0.
Confusion Matrix.
Collected over the test split and used to compute TP/FP/FN counts for each class.
C.6 Fairness and Consistency
To ensure a fair comparison with CADAR and VLM baselines:
  • All models receive the same raw input videos.
  • No prompts, symbolic structures, or additional annotations are provided to video baselines.
  • Clip preprocessing (20 frames, resizing, normalization) is identical across all supervised baselines.
Overall, these settings yield strong and transparent supervised video baselines, allowing us to isolate the contributions of CADAR’s neurosymbolic reasoning.
C.7 Model Architectures and Training
For our supervised-learning baselines—3D ResNet and ViViT. Training hyper-parameters are shown below. We conducted all training and evaluation using Google Colab with NVIDIA A100 GPUs (40GB memory). The experiments were run in a Linux environment via Colab’s remote infrastructure, conducted using PyTorch 2.6.0 with Python 3.11.
@@ C.7 Model Architectures and Training
We use ViViT-B/16x2 (pretrained on Kinetics-400) as the backbone for our ViViT model, and a 3D ResNet-18 for the ResNet baseline. Both models are trained on the full training set using inverse-frequency sampling to balance class distributions. The classification task involves five classes: the four attack types, along with an additional No Attack category. All input videos are resized to 400×400 pixels and normalized using a mean of [0.45, 0.45, 0.45] and a standard deviation of [0.225, 0.225, 0.225]. Each input consists of 20 uniformly sampled frames from a video clip. We use Focal Loss with a gamma of 2.0 as the training objective. Mixed precision (FP16) training is enabled via HuggingFace Accelerate for both models. ViViT is trained with a cosine learning rate scheduler, while the 3D ResNet uses a StepLR scheduler (step size = 10, gamma = 0.1). Final hyperparameter configurations for both models were selected based on the settings that yielded the highest validation accuracy.
Appendix D Vision–Language Model (VLM) Baseline Implementation Details
We evaluate three state-of-the-art multimodal VLMs as baselines: GPT-5-mini (OpenAI), Gemini-2.5-flash (Google), and Grok-4-fast (xAI). All three models support multi-image inputs and natural language reasoning, enabling them to perform frame-level attack classification without any training or fine-tuning. This section details our prompt design, image preprocessing strategy, and inference procedure, following our released implementation.
D.1 Input Frames and Preprocessing
Each AR video is converted into multiple reference and attack frames using the procedure in our VLM baseline code.We extract:
  • N reference frames uniformly from the first 30–40% of the video, and
  • M attack frames uniformly from the last 25%.
Images are resized to a maximum resolution of $800\times 800$ while preserving aspect ratio. For GPT-5 and Grok, images are encoded as JPEG base64 strings. Gemini receives PIL image objects directly, as required by the generate_content interface.
D.2 Unified Prompt for All VLM Baselines
To ensure fair comparison across models, all VLMs use the exact same prompt (text only; no model-specific tuning). We introduce the four base attack types used in CADAR—text modification, visual modification, obstruction, and injection—together with detailed natural-language definitions. The full prompt is shown below.
You are an AR video attack analyst. Your job is to identify the presence of 4 different kinds of AR attacks and non-attacks in the video with the scene context. The definitions for each attack are as follows:
@@ D.2 Unified Prompt for All VLM Baselines
A text modification attack manipulates textual information embedded in the environment, thereby changing the semantic message conveyed by the scene. For example, replacing a “NO PARKING FIRE LANE” sign with “FREE PARKING FIRE LANE” directly reverses the intended directive and can create safety hazards. Such attacks are consequential when text provides primary contextual or navigational cues.
@@
A visual modification attack alters an object’s appearance—its color, shape, or iconography—or repositions it implausibly. Examples include transforming a pedestrian-crossing sign into a bicycle-crossing sign, switching a green traffic signal to red, or relocating a stop sign to the center of an intersection. These manipulations can be quantified by analysing pixel-level discrepancies and deviations from expected spatial distributions.
@@
An obstruction attack hides or deletes critical objects, thereby eliminating essential information and distorting scene semantics. For example, erasing an emergency-exit sign in a fire evacuation scenario removes a safety-critical cue and disrupts the relational structure of the environment.
@@
An Injection attack introduces new, contextually inconsistent, or misleading objects that distort scene semantics. A fake detour sign placed on a clear street, or fictitious warning icons inserted into an AR overlay, may mislead human users and perception systems. Detection relies on identifying unexpected objects that conflict with the scene’s established structure and context.
A non-attack refers to a video in which the AR content remains semantically consistent with the physical environment and does not alter the user’s understanding of the scene.
@@
The following input will contain video frames and context. Please output the class. Note that there may be more than one kind of attack present in the video, or no attack at all. Along with the attack, mention the object the attack was performed on, or briefly describe the attack in the context of the video.
This prompt is prepended to the multimodal input for all VLMs.
D.3 Inference Procedure
Each VLM receives:
  • the textual prompt above,
  • scene context,
  • $N=20$ reference images, and
  • $M=5$ attack image at a time.
The models are called with their native multimodal APIs:
  • GPT-5-mini: OpenAI Chat Completions (image URLs encoded as base64).
  • Gemini-2.5-flash: generate_content with a mixed list of text + images.
  • Grok-4-fast: xAI Chat Completions (OpenAI-compatible JSON format with image URLs).
Each response is required to be valid JSON, following the schema below:
    "object": "<object>",
    "attack_type": "<classification>",
    "notes": "<short explanation>"
We apply a lightweight parser that extracts JSON blocks from the raw VLM outputs. Each model produces one prediction per attack frame. This mirrors video-level behavior and reduces variance inherent in single-frame VLM outputs.
D.4 Fairness Measures
To ensure a strict apples-to-apples comparison:
  • All VLMs use the same prompt with context, the same set of reference frames, and the same attack frames.
  • No model-specific prompt engineering, priming, or handcrafted reasoning chains are used.
  • All text is fed to each VLM exactly once per attack frame; no iterative refinement or self-consistency sampling is applied.
Because VLMs are not adapted to CADAR’s symbolic representation or temporal structure, these baselines reflect the true performance of general-purpose open-world vision–language models on cognitive attack detection.